OpenStudy (anonymous):

The half-equivalence point of a titration occurs half way to the end point, where half of the analyte has reacted to form its conjugate, and the other half still remains unreacted. If 0.300 moles of a monoprotic weak acid (Ka = 6.3 × 10-5 M) is titrated with NaOH, what is the pH of the solution at the half-equivalence point?

OpenStudy (ookawaiioo):

Use the henderson-hasselbach reaction

OpenStudy (ookawaiioo):

pH = pKa

OpenStudy (ookawaiioo):

pKa = -log(6.3x10^-5)

OpenStudy (ookawaiioo):

pKa = 5-log(6.3) uhhhhh...... 5 - Log2+Log3 ummmm.....5 - (0.3)+(.48) ooooooo......5 - (0.78) pH = (roughly) 4.22